OpenStudy (anonymous):

Write the standard form of the equation of the parabola that has the indicated vertex and whose graph passes through the given point. Vertex: (2, −1); point: (5, 17)

OpenStudy (nurali):

The formula for a parabola in vertex form is f(x) = a(x-h)² + k Since the vertex is (2,-1), plug those values in for h and k, but not for "a". y = a(x-2)²-1 Now plug (5,17) into this for x and y, so you can solve for "a". 17 = a(5-2)²-1 17 = 9a-1 18 = 9a a = 2 y = 2(x-2)²-1 <==ANSWER

OpenStudy (anonymous):

thank you

OpenStudy (nurali):

Anytime.